4(2-x)=21 sove for x please :)
Answer by wilft1(217)

ok, solve for x...
4(2-x) = 21
eliminate the parenthesis by multiplying the insides by 4
8-4x = 21
subtract 8 on both sides
-4x = 13
divide both sides by -4
x = - 13/4
or
x = -3.25